### v2

# BASED: Bundle-Adjusting Surgical Endoscopic Dynamic Video Reconstruction using Neural Radiance Fields

Abstract
Reconstruction of deformable scenes from endoscopic videos is important for many applications such as intraoperative navigation, surgical visual perception, and robotic surgery. It is a foundational requirement for realizing autonomous robotic interventions for minimally invasive surgery. However, previous approaches in this domain have been limited by their modular nature and are confined to specific camera and scene settings. Our work adopts the Neural Radiance Fields (NeRF) approach to learning 3D implicit representations of scenes that are both dynamic and deformable over time, and furthermore with unknown camera poses. We demonstrate this approach on endoscopic surgical scenes from robotic surgery. This work removes the constraints of known camera poses and overcomes the drawbacks of the state-of-the-art unstructured dynamic scene reconstruction technique, which relies on the static part of the scene for accurate reconstruction. Through several experimental datasets, we demonstrate the versatility of our proposed model to adapt to diverse camera and scene settings, and show its promise for both current and future robotic surgical systems. Project Website: https://sites.google.com/ucsd.edu/based-arclab-2023/.
Index Terms:
Endoscopic video, surgical scene reconstruction, bundle-adjusting reconstruction and pose estimation, implicit neural representation, deformable scene.
[…]
=-15pt
I INTRODUCTION
Surgical scene reconstruction from endoscopic videos is an important and challenging task in robotic-assisted minimally invasive surgery [1]. It also has widespread use in many downstream clinical applications such as surgical workflow analysis, image-guided robotic surgery, automation, surgical environment simulation, and augmented reality to train surgeons to be familiar with an operative environment and reconstruction of tissues and organs for preoperative assessment and planning. Many learning and automation algorithms for robotic surgery are reliant on accurate 3D reconstruction of the surgical scenes [2, 3, 4].
The majority of current approaches for reconstructing surgical scenes from endoscopic videos adopt a modular sequence of tasks, each tackled by independent models [5, 6, 7, 8, 9, 10, 11, 12, 13]. This fragmented methodology is susceptible to accumulating errors and lacks efficient task coordination. For instance, a significant number of existing methods either estimate or rely on ground truth camera poses before proceeding with scene reconstruction. Errors in pose estimation often lead to faulty scene reconstructions. Camera poses obtained through structure from motion (SfM) or simultaneous localization and mapping (SLAM) [14, 15, 16, 17, 18] often exhibit noise. When the camera is fixed to the end effector of a robotic arm, camera pose estimation can be accomplished through forward kinematics. However, these estimated poses are prone to errors stemming from inaccuracies in joint angle measurements attributed to instrument sag, cable stretch, positioning biases, and reading drift [19]. In reality, a further complication is that newer robotic systems are flexible, and these continuum robots (e.g., J&J Monarch Platform [20], Endomaster EASE System [21]) have complex mechanics that result in poor distal-camera localization. Furthermore, in addition to being modular, these methods will not generalize from static to deformable scene settings.
Ground Truth Color (left) and Depth (right)
RoDynRF (SOTA) color and depth
Refer to caption BASED (Ours) color and depth
Figure 1: BASED is a novel NeRF-based method that can be used in dynamic and deformable scenes with unknown camera poses. It can produce novel viewpoint renderings with robust color (left) and depth (right) reconstructions, even from monocular untracked camera images. Comparisons with state-of-the-art methods (RoDynRF shown above, EndoNeRF failed completely) show that it has leading performance in scene reconstruction. Depth in the ground truth row is the reference depth, estimated from [6].
Neural Radiance Fields (NeRF) [22] has shown tremendous performance in the domain of 3D reconstruction. They have been applied recently to stereo-endoscopic videos for deformable surgery scene reconstruction [23, 24]. However, some constraints in these methods are that the camera should be static with stereo camera information and be in a known position. Alternatively, some works have extended NeRF with camera pose optimization (e.g., [25, 26]). However, these methods generally assume the scene to be static.
More recently, 3D Gaussian Splatting (3DGS) [27] was proposed as an alternative of NeRF. 3DGS represents the sence with a cloud of 3D Gaussians and estimate the Gaussian parameters based on a efficient rendering algorithm. Similar to NeRF, this technique has been applied to deformable surgical scenes with fixed or known camera poses [28, 29], as well as simultaneous estimation of static scene representations and unknown camera poses [30]. However, although 3DGS has achieved state-of-the-art visual rendering quality images, its scene representation – the 3D Gaussians – does not capture the 3D geometric information of the scene accurately enough. This is because surgical videos are typically highly under-constrained; that is, the camera movement is relatively small and the videos are captured from similar perspectives. Moreover, since training of 3DGS primarily focuses on minimizing the difference between rendered and actual images rather than learning accurate 3D geometry of the scene, Gaussians with realistic colors but inaccurate positions relative to the camera can still achieve good rendering results in under-constrained scenes. Thus, the geometric accuracy of the Gaussians, specifically their positions, highly depends on the quality of the point cloud and camera poses initialized using off-the-shelf SfM methods like COLMAP [31, 32], which still struggle to handle certain surgical videos. Therefore, we develop our method based on NeRF and demonstrate its improved ability to capture scene geometry through several experiments.
In summary, we present the first NeRF-based method that can reconstruct deformable surgical scenes captured from untracked monocular endoscopic videos. The overall method combines a learnable pose layer that parameterizes camera poses, a deformation module that learns the deformation of a 3D point at a given time step with respect to a canonical position and a rendering module that takes the canonical coordinates of a 3D point at every time step along with the 2D camera directions and outputs the volumetric density and color information. We can solve for surgical scene representation learning, capturing more accurate geometric information without compromising on rendering capability for deforming scenes, even when the camera is a monocular camera, and in instances where the camera motion is unknown. This makes the method applicable to a wide range of robotic systems and semi-autonomous applications. The major contributions of this paper are thus:
  • A NeRF approach for dynamic and deformable scenes from data captured at unknown camera poses,
  • A method for describing multi-view correspondence loss, specifically catered to a dynamic and deformable setting, that enables establishing correspondence across spatiotemporal drift in tissues, and
  • A framework that offers a generic, unrestricted, and versatile solution to robotic surgical perception that is validated on several surgical scene reconstruction datasets.
II Related Work
II-A Traditional Scene Reconstruction Techniques
Structure from Motion (SfM) utilizes a series of two-dimensional images to construct a three-dimensional model of a scene or an object, and it is widely used in the domain of scene/organ/tissue reconstruction with endoscopic images [8, 9, 10, 11, 12, 13]. However, SfM techniques like COLMAP [31, 32] often fail to estimate camera poses for the entire length of sequences and often change the sequence of images. Simultaneous localization and mapping (SLAM) is a technique used to build a map of an unknown environment while simultaneously estimating the camera’s position and orientation in the environment. SLAM-based approaches [17, 16, 14, 15, 18] have been employed in the field of endoscopic scene reconstructions to model sparse and dense tissue surfaces. The scene is usually assumed to be static, and the estimated deformations do not always look natural.
II-B Static and Dynamic NeRF
NeRFs are implicit volumetric representations that encode the appearance and geometry of a 3D scene. It learns a continuous representation of the scene, enabling novel view renderings, unlike SLAM/SfM techniques that usually learn a discrete representation. It takes a 3D point coordinate and a 2D camera viewing direction and outputs RGB color information and the density of the point. The original NeRF model strictly catered to static scenes. D-NeRF [33] is the first work that expanded the NeRF model to learn deformable scenes. It consists of two separate models that progressively learn the deformation and the volumetric scene reconstruction simultaneously. Robust Dynamic NeRF, or RoDynRF [34], tries to reconstruct dynamic scenes by estimating the camera poses using only the static part of the scene, which is not suitable for surgical scene reconstruction given that the scenes are highly deformable with a very low degree of rigidity. EndoNeRF [23] builds on D-NeRF for endoscopic scenes by adding a tool-guided ray casting and by sampling points along a ray closer to the tissue surfaces using a Gaussian transfer function but relying on stereo depth. EndoSurf [24] further improved upon EndoNeRF [23] by employing three networks to learn the deformation, signed distance function (SDF), and color of a viewpoint. However, as seen in Figure 1, these methods strictly cater to cases where the camera is static. If extended to moving camera sequences with camera poses estimated by COLMAP, the rendering quality degrades significantly.
II-C Static and Dynamic Gaussian Splatting
3DGS is a fast, learnable rendering technique for novel view synthesis. It uses 3D Gaussians for scene representation and a novel tile-based rasterizer for fast 3D Gaussian rendering. Given a set of camera poses and their corresponding 2D views, a cloud of 3D Gaussian representation (either randomly initialized or sparsely initialized via structure from motion point clouds) can be learned end-to-end. It yields state-of-the-art image quality as well as real-time inference speed. While the original work focused on static scenes, dynamic and deformable scenes have also been explored in recent months.
Dynamic 3D Gaussians [35] is the first to utilize 3D Gaussian splatting for multiview videos. It builds a static scene using 3DGS for the first frame, then incrementally learns the offset for each 3D Gaussian. However, the content is limited to the initial frame. Hence, it cannot model novel objects appearing in the middle of the video. More recent 3DGStream [36] has improved the efficiency of offset cache storage and enabled new Gaussian spawning. However, these models can only work with multiview images and rely heavily on the information from the first frame.
Yang et al. [37] and Wu et al. [38] both use a deformation field to model offsets to a canonical space of explicit 3D Gaussians, whereas [39] has increased GS dimension into 4D. However, these methods assume ground truth camera poses are known. Our model does not have this assumption.
II-D Joint Pose Estimation and Scene Reconstruction
Bundle adjusting neural radiance fields (BARF) [40] first proposed an end-to-end NeRF-based framework that can jointly estimate camera extrinsics and reconstruct the 3D scene. It implements a coarse-to-fine technique to gradually activate the higher frequency components of the positional encoding scheme that aids in pose refinement, and only uses the photometric loss for backpropagation through the network. There have been numerous follow-up works ( [26, 41, 42, 43, 44, 25, 45, 46, 47]) that improved BARF by optimizing its training time, adding prior losses and making it robust to in-the-wild sequences. More recently, similar ideas have also been explore with 3DGS [30, 48]. However, all of these methods specifically deal with static scenes. RoDynRF  is a recent paper, to the best of our knowledge, that has first implemented estimating camera poses while simultaneously reconstructing dynamic scenes. They combine the pose estimation module from BARF with the radiance field module of Dynamic NeRF [49]. Although RoDynRF’s problem formulation is similar to ours, it cannot be easily adapted in surgical scene settings, requiring the dynamic pixels to be masked out prior to training, where a motion mask is generated using Mask R-CNN [50] and optical flow [51]. This is hard to acquire as the entire surgical scene may be deformable and textureless. Hence, as seen in Figure 1, it cannot be generalized for endoscopic videos. Our proposed model overcomes the limitations of EndoNeRF and EndoSurf by expanding to moving cameras with unknown poses. It does not constrain part of the scene to be static for it to be able to learn deformations, thus also overcoming the limitations of RoDynRF.
[…]
III Methodology
Figure 2 gives a high-level overview of our method. The overall model consists of three parts: (1) the pose estimation module (see III-A), (2) the deformable module (see III-B), and (3) a canonical NeRF module (see III-C).
III-A Camera Pose Module
We use a camera pose module to model the appearance of the rigid-body motion of the scene either caused by a moving camera or moving scene as is shown in Figure 2 (A). We follow the convention of previous work [40] and model the camera pose with $P_{i}=[R_{i}|t_{i}]\in SE(3)$ representing camera-to-world transform of camera $i$, where $R_{i}\in SO(3)$ denotes the rotation and $t_{i}\in\mathbb{R}^{3}$ denotes the translation. We use $K\in\mathbb{R}^{3\times 3}$ to represent the intrinsic matrix.
In terms of implementation details, the poses are parameterized as a learnable parameter $P_{i}\in\mathbb{R}^{9}$ for $i\in[1,N]$ for $N$ frames. The first six columns of the matrix refer to the first two rows of the camera rotation matrix, and the last three values refer to the camera translation vector. It should be noted that one can estimate the third row of the rotation matrix by taking a cross-product of the other two rows. The pose layer is initialized to identity and updated for the first 200 iterations, after which the layer is frozen.
III-B Deformation Module
After the camera pose module explains away the rigid-body motion in the image, the deformation module will model the non-rigid-body deformation of the scene from a canonical state at $t_{0}$ as is shown in Figure 2 (B). The deformation module $\psi_{\Delta}(x,t)$ takes as input 3D spatial coordinates $x\in\mathbb{R}^{3}$ and a time step $t$ and learns the deformation $\Delta x$ of the coordinate $x$ from time step zero:
$\psi_{\Delta}(x,t)=\begin{cases}\Delta x&\text{if }t\neq 0\\ 0&\text{if }t=0\end{cases}$ (1)
The idea is that the deformation can help us learn the position estimate of a point at the $0^{th}$ time interval, which will be assumed the canonical configuration. If $x_{0}$ is the coordinate of a point at time $t=0$ and it deforms to $x_{t}$ at time t, then we can say that $\psi_{\Delta}(x_{t},t)+x_{t}=x_{0}$.
III-C Canonical NeRF Module
This module will capture the implicit 3D representation of the canonical state of the scene. The canonical module $\psi_{x}$ can be assumed to be a vanilla NeRF model, which takes as input a spatial 3D location and 2D camera orientations, and outputs color (RGB values) and the volumetric density of the location. $x_{0}$ calculated above is the input to the canonical model along with camera viewing directions $d$. The rendered output of the canonical model (color $c$, and density $\sigma$) is corresponding to the deformed point $x_{t}$ at time $t$.
$\psi_{x}(x_{t}+\psi_{\Delta}(x_{t},t),d)=c,\sigma$ (2)
We apply conventional positional encoding [33] to all the inputs, $<(sin(2^{l}\pi p),cos(2^{l}\pi p))>_{0}^{L}$, with $L=10$ for 3D point coordinates $x$ and time frame $t$, and $L=4$ for camera directions $d$ For more information about volumetric rendering, please refer to the original NeRF paper [22].
III-D Tool mask-guided ray casting and Stereo Depth Cue Ray Marching
Surgical tools occlude part of the scene in most of our sequences. Instead of shooting rays randomly all through the image, we adopt a tool mask-guided ray casting strategy [23]. Using binary tool masks, an importance map is created which highlights the regions containing tool pixels. During training, instead of random ray sampling throughout the image, we only sample pixels from this distribution using inverse transform sampling. Secondly, we sample pixels close to the tissue surface using a Gaussian transfer function and use a depth refinement strategy to get rid of corrupt artifacts in the estimated depths.
III-E Losses
Figure 3: Overview of the correspondence loss: Matching pixels are chosen from a random pair of images (having correspondence above a certain threshold) using PDC-Net [52]. The matching pixels are projected from 2D pixel space to their 3D positions in the world frame, and then further passed through the deformation model to get their canonical 3D positions. The difference between the mapped 3D points is used to calculate the flow correspondence loss. The loss is designed to pull the outputs of correspondence points returned by the deformation model closer together towards the same canonical point denoted as $x_{0}$.
Our model uses photometric loss, correspondence loss, and estimated depth Loss for monocular dynamic camera deformable scene reconstruction.
RGB Photometric Loss: We use photometric loss based on conventional setup in NeRF models [22]. At every training iteration, we compare the rendered color of the sampled pixels with the ground truth colors using the $L_{2}$ loss, and this loss is backpropagated through the entire network with all three modules.
$L_{pho}=\frac{1}{N}\Sigma_{i=1}^{N}||\hat{C_{i}}(p,t)-C_{i}(p,t)||_{2}^{2}$ (3)
where $\hat{C_{i}}$ is the rendered color and $C_{i}$ is the ground truth color of the pixel $p$ at time $t$ with respect to frame $i,~{}i=0...N$.
Dynamic Multi-View Correspondence Loss: We introduce a novel 3D multi-view correspondence loss for dynamic scenes to offer extra constraints on pose estimation. Figure 3 gives a high-level idea of this loss. For a given pair of images, $I_{i}^{a}$ (viewpoint $P_{i}$ at time $t_{a}$) and $I_{j}^{b}$ (viewpoint $P_{j}$ at time $t_{b}$) with matching pixels $p^{a}\in I_{i}^{a}$ and $q^{b}\in I_{j}^{b}$, both pixels $p^{a}$ and $q^{b}$ should project to the same point in 3D space in the canonical space. $z_{p}^{a}$ and $z_{q}^{b}$ are the estimated depths of the pixels $p^{a}$ and $q^{b}$ respectively. Let $f_{P}^{-1}$ be the function that maps the 3D coordinate of a pixel given its estimated depth and its respective camera viewpoint. Let $x_{p}^{a}\in\mathbb{R}^{3}$ be the 3D coordinate of pixel $p^{a}$ at time $t_{a}$ and $x_{q}^{b}\in\mathbb{R}^{3}$ be the 3D coordinate of pixel $q^{b}$ at time $t_{b}$. Then, the estimated coordinates are
$x_{p}^{a}=f_{P_{i}}^{-1}(p^{a},z_{p}^{a})$ (4)
$x_{q}^{b}=f_{P_{j}}^{-1}(q^{b},z_{q}^{b})$ (5)
If there was no deformation, $x_{p}^{a}$ and $x_{q}^{b}$ should be the same point in 3D space. However, since this is a deformable scene, we need to map both the points back to their canonical state in order to calculate the loss:
$\hat{x}_{p}^{a}=\psi_{\Delta}(x_{p}^{a},t_{a})+x_{p}^{a}$ (6)
$\hat{x}_{q}^{b}=\psi_{\Delta}(x_{q}^{b},t_{a})+x_{q}^{b}$ (7)
Then, the loss becomes:
$L_{corr}=w(L_{\delta}(\hat{x}_{p}^{a},\hat{x}_{q}^{b}))$ (8)
where $w$ is loss weight term and $L_{\delta}$ is the Huber loss function.
Estimated Depth Loss: In spite of only having 2D images, we can use an estimated depth loss to further assist better reconstruction when the camera pose estimation is optimized. The depth estimates can be found via an off-the-shelf image-to-depth estimator. This loss is only used to optimize the deformable and canonical modules after the pose layer is frozen. Applying the estimated depth loss before the camera pose matrix is optimized leads to abrupt artifacts. We collect depths estimated from RGB images via existing pre-trained methods [53][6] during training. We compare the rendered depth values of the sampled pixels with the corresponding estimated depth values using a $L_{2}$ loss. The gradient from this loss is only backpropagated through the deformable NerRF modules, after the pose layer is frozen.
$L_{d}=\frac{1}{N}\Sigma_{i=1}^{N}||\hat{D_{i}}(p,t)-D_{i}(p,t)||_{2}^{2}$ (9)
where $\hat{D_{i}}(p,t)=\frac{1}{\sigma_{i}(p,t)}$ is the rendered depth, $\sigma_{i}$ is the rendered density and $D_{i}(p,t)$ is the reference depth of the pixel $p$ at time $t$ with respect to frame $i\in[1,N]$ for $N$ frames. Note that if a hardware solution to depth is available (e.g., stereo cameras or RGB-D sensor, the ground truth depth can be used as well to calculate an estimated depth loss.
IV Experiments
IV-A Implementation Details
All of the experiments have been performed on a single Nvidia RTX-A6000 GPU. The weights of the NeRF network and the pose estimation network are simultaneously updated with a learning rate of 5e-4. After 200 iterations, the pose layer is fixed, and only the NeRF model is updated for another 100K iterations. In the first stage, one batch consists of all of the training images, because it requires more views of the scene for better pose estimation. In the second stage, one batch consists of a single training image for faster training. Tool masks are manually created for the Hamlyn dataset sequences, partly using Segment Anything Model [54]. We use an off-the-shelf dense pixel correspondence network, PDCNet [55], to obtain corresponding matching pixels between a pair of images for calculating the multi-view correspondence loss $L_{corr}$.
TABLE I: Ablation studies on Hamlyn Rectified18-1 dataset with dynamic motions.
[…]
$\checkmark$ indicates loss usage.
BA stands for bundle-adjusting.
EndoNeRF experiment is using COLMAP for camera pose estimation.
Ground Truth BASED w/o $L_{d}L_{corr}$ BASED w/o $L_{d}$ BASED w/o $L_{corr}$ BASED (final)
Figure 4: Ablation analysis on Hamlyn Rectified 18-1 dataset shows how different losses bring about an improvement in the rendered results from the final model. Please note that the depth in the Ground Truth column is the reference depth.
IV-B Datasets
We evaluated BASED on the following datasets:
Hamlyn Dataset [56]: We choose three videos from the Hamlyn dataset of robotic surgery with deformable scenes. Hamlyn dataset contains weak textures, and the scene is often occluded by blood and surgical instruments. Rectified04 contains 158 frames, and it appears to have slight camera motion. The camera in Rectified04 is mostly static, whereas Rectified18 exhibits slightly more camera motion, so two subsets of Rectified18 are chosen from the whole video with 113 frames in each. The frames are sampled at regular intervals (every 10 frames), as the original videos were too large. 10 percent of the total images were used for testing. We use [6] to generate reference depth needed for calculating the estimated depth loss.
EndoNeRF da Vinci dataset [23]: We used only two datasets used in the EndoNeRF paper that are publicly available. They last for 4-8 sec with 15 fps. Each of the videos is taken from a single static viewpoint setting. One of the videos demonstrates soft tissues being drastically pulled in 2 seconds, while the other video demonstrates soft tissue cutting with challenging topological changes. As mentioned in the original paper, the depth maps are obtained using STTR-LIGHT [53] from stereo images. The train-test split is the same as it is used in the original paper.
IV-C Metrics
The quality of image rendering is evaluated through commonly used metrics, including (1) Peak Signal-to-Noise Ratio (PSNR), (2) Learned Perceptual Image Patch Similarity (LPIPS) [57], and (3) Structural Similarity Index (SSIM) [57]. We also quantitatively evaluate the 3D reconstruction performance by comparing predicted depth maps with the reference depth (refer to Section IV-B for details), using standard metrics for depth evaluation [58]. With this, we are treating the reference as pseudo ground truth. It should be noted that while the quality this pseudo ground truth may be slightly inferior than that of true depth, using it is still a common strategy in endoscopic video analysis as true ground truth can be hard to obtain in a surgical environment [59].
IV-D Results
TABLE II: Quantitative analysis of BASED and State-of-the-Art on the Hamlyn Dataset.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ EndoNeRF RoDynRF EndoGS Ours EndoNeRF RoDynRF EndoGS Ours EndoNeRF RoDynRF EndoGS Ours Rectified04-1 (scds) - 14.902 - 20.947 - 0.468 - 0.621 - 0.577 - 0.269 Rectified18-1 (mcds) 22.838 25.700 30.821 32.277 0.817 0.874 0.928 0.944 0.273 0.192 0.100 0.106 Rectified18-2 (mcds) 29.572 24.690 35.969 33.017 0.869 0.875 0.956 0.936 0.235 0.201 0.079 0.103
‘-’ denotes that the model fails to complete the task.
Best results are bolded; second best are underlined.
TABLE III: Quantitative analysis of BASED and EndoGS depth estimation on the Hamlyn Rectified18-2 dataset.
[…]
Figure 5: Camera poses and point cloud visualization of (a) EndoGS and (b) BASED on hamlyn dataset Rectified18-2. The colors assigned to the cameras correspond to their timestamps in the video.
IV-D1 Ablation Study
Table I and Figure 4 provide both quantitative and qualitative analysis of the effect of the various losses on the performance of our model. As shown in Table I, bundle-adjusting leads to significant improvement in image rendering quality compared with EndoNeRF [23], a state-of-the-art NeRF-based model developed for a similar task. Also, training with estimated depth guidance $L_{d}$ and correspondence loss $L_{corr}$ improves rendering quality, respectively. Figure 4 shows that estimated depth guidance clearly improves the reconstructed textures and depths. Additionally, correspondence loss $L_{corr}$ leads to better results, as shown quantitatively by image rendering results and qualitatively by predicted depth quality, indicating that the correspondence loss may have improved the estimated camera poses by instilling multi-view consistency in this dynamic deformable scene, which in turn leads to better convergence of the deformable NeRF model. However, the sequences are still severely underconstrained (camera movement/viewing angles of the scene are still limited in those sequences, which is common in surgical scenes), hence the correspondence loss only showed a relatively slight improvement in depth reconstruction. The tool masks for sequence - Rectified04-1 from the Hamlyn dataset does not always completely mask out the tools, which is why the resultant rendered images partially have tools in them (Figure 1), and not a transparent overlay like we see for the other rendered results.
Ground Truth BASED (Ours) EndoNeRF RoDynRF         EndoGS
Figure 6: Qualitative analysis of our proposed model BASED compared with RoDynRF, EndoNeRF and EndoGS (with COLMAP estimated poses) for sequences obtained from the public Hamlyn dataset. Please note that the depth in the Ground Truth column is the reference depth.
IV-D2 Comparison on Hamlyn Dataset
The main dataset we are evaluating is the Hamlyn dataset. We compare BASED with state-of-the-art methods including EndoNeRF, RoDynRF and EndoGS, for deformable scene reconstruction with obvious camera motions. The image rendering quality are quantitatively and qualitatively compared in Table II and Figure 6, respectively. The 3D scene reconstruction performance is evaluated through depth maps predicted by the models, which are also provided in Figure 6. Our model outperforms both EndoNeRF and RoDynRF by a significant margin in both image rendering and 3D reconstruction.
TABLE IV: Quantitative analysis of BASED and state-of-the-art on the EndoNeRF da Vinci dataset.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ EndoNeRF RoDynRF EndoGS Ours EndoNeRF RoDynRF EndoGS Ours EndoNeRF RoDynRF EndoGS Ours pulling soft tissues 35.832 27.502 36.750 35.097 0.941 0.900 0.925 0.936 0.058 0.158 0.135 0.059 cutting tissues twice 35.720 21.105 37.372 35.706 0.935 0.700 0.964 0.934 0.063 0.280 0.050 0.073
Best results are bolded; second best are underlined.
TABLE V: Quantitative analysis of BASED and EndoGS depth estimation on the EndoNeRF ”Cutting Tissues Twice” dataset.
[…]
On the other hand, BASED is still competitive in image rendering compared with EndoGS, which is based on 3DGS. 3DGS has inherent advantages in image rendering, as the color and position of the Gaussians can be easily updated to fit the camera view. However, the Gaussian cloud estimated by EndoGS may not accurately represent the true geometric shape of the environment, which is critical for many downstream tasks in surgical applications. To demonstrate the superiority of our method in capturing the 3D geometric information of the scene, beyond Figure 6, we also quantitatively measure the quality of predicted depth based on the reference depth, as shown in Table III. BASED provides better depth maps than EndoGS both quantitatively and qualitatively. Furthermore, we visualize the camera pose along with the point cloud estimated by EndoGS and BASED in Figure 5. Given the small camera motion in this dataset, the true camera poses are expected to be close over time. BASED aligns more closely with this pattern compared to EndoGS. A major drawback of EndoNeRF and EndoGS is that they have only been tested on video sequences with single-view static cameras, which are severely under-constrained. To apply EndoNeRF and EndoGS to dynamic scenes, both models use COLMAP to estimate camera poses. Thus, these two models fail when COLMAP fails to predict camera poses as seen for Rectified 04-1 in Table II. This clearly demonstrates the superiority of our method. Also, even in cases where COLMAP is able to estimate camera poses, our model, BASED, more accurately captures the dynamic and deformable aspects of the scene, as shown by the quantitative and qualitative results provided.
Ground Truth BASED (Ours) EndoNeRF RoDynRF         EndoGS
Figure 7: Qualitative analysis of our proposed model BASED with RoDynRF and EndoNeRF for two publicly available sequences - ”Cutting Tissues Twice”[23] guided by stereo depth.
IV-D3 Comparison on EndoNeRF da Vinci Dataset
We further show that our model can generalize toward deformable surgical scenes with stationary cameras on the sequences used in the EndoNeRF paper: Cutting tissues twice and Pulling soft tissues. Table IV and Figure 7 show a quantitative and qualitative evaluation of the performance of our model versus baseline EndoNeRF, RoDynRF. We can see that our model is on par, if not better than EndoNeRF, and performs significantly better than RoDynRF. In Figure 8, our model is able to generalize better to novel view renderings compared to EndoNeRF which has been strictly trained with static identity camera poses. Hence, our model retains EndoNeRF’s capacity for static deformable monocular videos.
Furthermore, we still compare our model with EndoGS on static surgical scenes. From Table V and Figure 7, both qualitative and quantitative results show although EndoGS shows its inherent advantage in image rendering, it still almost fails in geometric shape capturing even on static camera scenes, which is very crucial for real-world applications of 3D scene reconstruction.
EndoNeRF               BASED
Figure 8: Qualitative analysis of our proposed model BASED and EndoNeRF on ”Cutting Tissues Twice” dataset for novel view renderings. Highlighted sections show how BASED is able to avoid different artifacts in the images, resulting in a sharper image.
V Conclusion
We propose BASED, a network that simultaneously estimates dynamic scene motions and reconstructs deformable surgical scenes from monocular videos. To the best of our knowledge, this is the first work that tries to reconstruct scenes from video sequences where the static and the dynamic parts cannot be clearly differentiated from one another, and also the first work that attempts to reconstruct deformable surgical scenes with unknown camera poses from monocular videos. Through our experiments, we have shown significant improvement over the existing state-of-the-art methods in this domain, and also potential for future research. Limitations of the presented model include slow optimization time (similar to EndoNeRF at several hours), though many recent strategies are drastically cutting this time by orders of magnitude (including ours [26]). Some potential future directions could be removing the dependency of BASED on reference depths, reducing optimization time, and automatic tool mask generation.